OpenStudy (anonymous):

How Do I Establish a Transformation Identity? y=M sin Bt + N cos Bt y = sqrt( M^2 + N^2) sin (Bt+C) C is any angle( in radians if t is real) having P=(M,N)on its terminal side [hint a first step is the following: M sin Bt + N cos BT = sqrt( M^2 + N^2)/sqrt( M^2 + N^2)*(M sin Bt + N cos Bt)

OpenStudy (campbell_st):

for a sin x + b cos x can be written as rsin(x + c) \[r = \sqrt{a^2 + b^2}\] tan c = b/a